Assume you invest $200 per month in your 401 K account in U.S. Treasury Bill, given your age (27years and 3 months), how much you will accumulate at age 67?...interest rate of 15.5%
Answer #1

Annuity= $200 per month

Present age= 27*12 + 3 = 327 months

Target age= 67*12= 804 months

N= 804-327 = 477

Rate= 15.5%/12

Amount accumulated= FV= Annuity*((1+r)^n-1)/r

= 200*((1+0.155/12)^477 -1)/(0.155/12))

=$7,040,256.88